carmelle:

Factor completely: 9x^4 + 12x^3 + 21x^2. 3(3x^4 + 4x^3 + 7x^2) 3x(3x^3 + 4x^2 + 7x) 3x^2(3x^2 + 4x + 7) Prime I think its A

Kamyoboi:

i think it iz da 3rd one

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Kamyoboi i think it iz da 3rd one \(\color{#0cbb34}{\text{End of Quote}}\) why. explanation-?

dude:

mm you want to find the common factor between \(x^4\), \(x^3\) and \(x^2\) Do you know what value it would be?

carmelle:

x^2

snowflake0531:

Then find the GCM of 9, 12, and 21

dude:

Yeah thats all you needed to really do (you'd need to solve for the factor for each coefficient if you weren't given similar options) 3(3x^4 + 4x^3 + 7x^2) 3x(3x^3 + 4x^2 + 7x) \(\color{orange}{3x^2}\)(3x^2 + 4x + 7) You can check it by expanding it \(\color{orange}{3x^2}(\color{blue}{3x^2} + \color{green}{4x} + \color{red}{7})\) \((\color{orange}{3x^2}\times\color{blue}{3x^2})=9x^4\) \((\color{orange}{3x^2}\times\color{green}{4x})=12x^3\) \((\color{orange}{3x^2}\times\color{red}{7})=21x^2\)

carmelle:

tysm